OpenStudy (anonymous):

Show all work in dividing y^2-3y+12/y^2+5y+6 by y^2+2y-8/y^2+y-12. Use complete sentences to explain how to simplify this expression. Please help!!!

OpenStudy (mertsj):

\[\frac{y^2-3y+12}{y^2+5y+6}\ \div\frac{y^2+2y-8}{y^2+y-12}\] \[\frac{y^2-3y+12}{(y+3)(y+2)}\times\frac{(y+4)(y-3)}{(y+4)(y-2)}=\frac{(y^2-3y+12)(y-3)}{(y+3)(y+2)(y-2)}\]

OpenStudy (mertsj):

I have a feeling that you posted this problem incorrectly. Most of these problems reduce more than this one does.

OpenStudy (anonymous):